OpenStudy (anonymous):

if f(x)=2x+3 and g(x)=x^2, then g[f(-3)]=?

OpenStudy (anonymous):

9

OpenStudy (zehanz):

First , calculate f(-3)=2*(-3)+3=... Put that result in g, so square it. That will be the answer.

OpenStudy (stamp):

\[f(x)=2x+3\]\[g(x)=x^2\]Find g(f(-3)).\[g(f(-3))=(f(-3))^2\]Find f(-3).\[f(-3)=2(-3)+3\]\[f(-3)=(-6)+3\]\[f(-3)=-3\]Substitute.\[g(f(-3))=g(-3)=(-3)^2\] g(f(-3)) = 9

OpenStudy (anonymous):

did you get the explanation above or should i type it?